# Intelligent Graybox Fuzzing via ATPG-Guided Seed Generation and Submodule Analysis

### v1

@@ I Introduction
In this work, we present PROFUZZ, a novel hardware fuzzing framework based on DGF, specifically designed to overcome the challenges, such as abstraction mismatches, limited coverage precision, and poor scalability. Unlike existing methods, PROFUZZ operates directly at the hardware’s native abstraction level, allowing it to accurately model inherent hardware behaviors and utilize coverage metrics that reflect hardware-specific characteristics. Additionally, our framework seamlessly integrates with industry-standard Electronic Design Automation (EDA) tools, facilitating easy adoption within established design and verification workflows. Our main contributions are summarized as follows:
@@ IV-C Evaluation Results
Table III presents the coverage values (in %) for different benchmarks with target submodule configuration and respective time (in seconds), demonstrating the ability of PROFUZZ to isolate and verify signal groups that span multiple modules and instances.
TABLE III: Evaluation for Generated Target Submodule
@@
TABLE IV: Comparing the performance of PROFUZZ with RFUZZ and DirectFuzz on various designs and highlighting the coverage improvement and speedup achieved by PROFUZZ over DirectFuzz
Benchmark
(Target Inst.)
#Signals
Target Inst.
Cell %
RFUZZ[12] DirectFuzz[25] PROFUZZ (This work)
@@
PROFUZZ achieves high efficiency in smaller datapath-centric designs like FIR, reaching 100% coverage with only 32.5% of internal signals targeted, completing in 4.81 seconds. The IIR design, with just 15.1% of nodes selected, also reaches full coverage in 4.31 seconds. Despite a moderate target ratio in GPIO (35.9%), PROFUZZ achieves 99.39% coverage in just 2.36 seconds. In more complex designs like ORCTRL, which feature larger input spaces and deeper logic, the framework still delivers 94.54% coverage with only 12.5% of internal nodes selected, showcasing its robustness in handling control-heavy modules. Cryptographic cores like DES3 and SHA256 achieve 98.62% and 98.09% coverage, respectively, in a short time, demonstrating strong observability despite their algorithmic complexity. The instrumentation overhead for large designs, such as PicoRV32, is effectively mitigated using the target submodule generation strategy, which enables directed seed generation focused on specific regions, as presented in Table III. Fig. 5 shows the coverage (%) and number of target signals for various benchmarks selected using two different cost functions: (1) nets with $\mathcal{C}_{i}\geq 0.3$ and (2) nets in the top 10 percentile ($P_{10}$) of the ranked cost function distribution. PROFUZZ achieves over 90% coverage across all benchmarks except SPI. The number of target signals is also highlighted to demonstrate PROFUZZ’s scalability. Notably, coverage remains consistent across both cost function configurations, demonstrating the framework’s ability to effectively cover target nets ranked by their cost function values.
@@ IV-D Comparison with RFUZZ and DirectFuzz
Table IV compares the performance of PROFUZZ with existing state-of-the-art techniques, including RFUZZ[12] and DirectFuzz[25]. As shown in previous subsections, PROFUZZ demonstrates superior efficacy by covering a significantly larger set of target signals compared to RFUZZ and DirectFuzz. While covering fewer target sites in smaller designs (e.g., UART) is easier, PROFUZZ performs better on more complex designs (e.g., Sodor5Stage). For a fair comparison, we used the same number of target signals as reported in DirectFuzz for each benchmark. As seen in Table IV, PROFUZZ achieves higher average coverage, with 11.66% improvement and a 2.76× speed-up over RFUZZ and DirectFuzz across designs of varying complexities. Notably, RFUZZ and DirectFuzz struggle with FFT, where the number of target sites (107) is the highest among all tested designs.
Fig. 6 compares the number of target sites handled by PROFUZZ and DirectFuzz across multiple benchmarks, showing a 2× to 30× increase in the number of target sites covered by PROFUZZ. This highlights the scalability and effectiveness of our approach, demonstrating its ability to handle a much larger number of target sites, particularly for complex hardware designs.

### v2

@@ I Introduction
In this work, we present PROFUZZ, a novel hardware fuzzing framework based on DGF, specifically designed to overcome the challenges, such as abstraction mismatches, limited coverage precision, and poor scalability. Unlike existing methods, PROFUZZ operates directly at the hardware’s native abstraction level, allowing it to accurately model inherent hardware behaviors and utilize coverage metrics that reflect hardware-specific characteristics. Additionally, our framework seamlessly integrates with industry-standard Electronic Design Automation (EDA) tools, facilitating easy adoption within established design and verification workflows.
Our main contributions are summarized as follows:
@@ IV-C Evaluation Results
Table III presents the coverage (in %) and the runtime (in seconds) for the evaluation benchmarks with randomly selected target signals. PROFUZZ demonstrates high efficiency on datapath-centric designs such as FIR and IIR, achieving 100% coverage while targeting only 32.5% and 15.1% of internal signals, respectively, in under 5 seconds. GPIO attains 99.39% coverage in 2.36 seconds despite a moderate target ratio (35.9%). For more complex control-heavy designs like ORCTRL, PROFUZZ achieves 94.54% coverage while targeting just 12.5% of internal nodes, demonstrating strong robustness. In contrast, designs such as PWM and SPI require over 50% of internal nodes to be targeted, achieving 53.24% and 75.63% coverage, respectively, reflecting the increased difficulty of covering dense target sets; nonetheless, SPI reaches 75.63% coverage in only 5.18 seconds. The UART design targets 64.6% of internal signals and achieves 93.81% coverage in 15.21 seconds, illustrating scalability to medium-complexity modules. Cryptographic cores (DES3, SHA256) achieve over 98% coverage within short runtimes, indicating strong observability despite algorithmic complexity. For I2C, with 1209 target nodes, the proposed framework achieves 94.10% coverage in approximately 240 seconds, further highlighting scalability of PROFUZZ. However, for large designs such as PicoRV32, DFT, and IDFT, full instrumentation incurs significant overhead ($\blacktriangle$) and resource usage (in Table III), limiting practicality. To address this, we employ a target submodule generation strategy, with results in Table IV, enabling scalable analysis of large-scale designs.
TABLE III: Evaluation for Random Selection of Target Signals
[…]
ORCTRL 415 94.54% 20.41
GPIO 325 99.39% 2.36
FIR 1049 100% 4.81
PWM 2008 53.24% 10.38
SPI 2857 75.63% 5.18
I2C 1209 94.10% 240.37
UART 983 93.81% 15.21
PicoRV32 5287 $\blacktriangle Ins Overhead\blacktriangle$
IIR 423 100% 4.31
DFT 9145 $\blacktriangle Ins Overhead\blacktriangle$
IDFT 9518 $\blacktriangle Ins Overhead\blacktriangle$
DES3 470 98.62 3.64
SHA256 993 98.09 10.09
[…]
$\nabla$
#Target Signals denotes the number of signals selected randomly from the design, ranging between 2%-50% of the total signals.
Table IV presents the coverage values (in %) for different benchmarks with target submodule configuration and respective time (in seconds), demonstrating the ability of PROFUZZ to isolate and verify signal groups that span multiple modules and instances. The instrumentation overhead for large designs, such as PicoRV32, is effectively mitigated using the target submodule generation strategy, which enables directed seed generation focused on specific regions, as presented in Table IV. For PicoRV32, with 455 target nodes, PROFUZZ achieves 60.42% coverage, while for ORCTRL it achieves similar coverage (with 415 nodes) under 20 seconds, but a more deeply embedded subset of 200 nodes requires nearly 300 seconds, highlighting the increased effort to activate nested logic. Smaller modules like GPIO, FIR, and I2C show consistent performance due to easier access to target sites. UART achieves 95.63% coverage for 219 target nodes, comparable to the full-module result of 93.81% with 983 targets (Table III), further validating the submodule approach.
TABLE IV: Evaluation for Generated Target Submodule
@@
Fig. 5 shows the coverage (%) and number of target signals for various benchmarks selected using two different cost functions: (1) nets with $\mathcal{C}_{i}\geq 0.3$ and (2) nets in the top 10 percentile ($P_{10}$) of the ranked cost function distribution. PROFUZZ achieves over 90% coverage across all benchmarks except SPI. The number of target signals is also highlighted to demonstrate PROFUZZ’s scalability. Notably, coverage remains consistent across both cost function configurations, demonstrating the framework’s ability to effectively cover target nets ranked by their cost function values.
TABLE V: Comparing the performance of PROFUZZ with RFUZZ and DirectFuzz on various designs and highlighting the coverage improvement and speedup achieved by PROFUZZ over DirectFuzz
Benchmark (Target Inst.) #Signals Target Inst. Cell % RFUZZ[12] DirectFuzz[25] PROFUZZ (This work)
@@ IV-D Comparison with RFUZZ and DirectFuzz
Table V compares the performance of PROFUZZ with existing state-of-the-art techniques, including RFUZZ[12] and DirectFuzz[25]. As shown in previous subsections, PROFUZZ demonstrates superior efficacy by covering a significantly larger set of target signals compared to RFUZZ and DirectFuzz. While covering fewer target sites in smaller designs (e.g., UART) is easier, PROFUZZ performs better on more complex designs (e.g., Sodor5Stage). For a fair comparison, we used the same number of target signals as reported in DirectFuzz for each benchmark. Notably, RFUZZ and DirectFuzz struggle with FFT, where the number of target sites (107) is the highest among all tested designs. As seen in Table V, PROFUZZ achieves higher average coverage, with 11.66% improvement and a 2.76× speed-up over RFUZZ and DirectFuzz across designs of varying complexities.
Figure 7: Coverage convergence of PROFUZZ with varying number of target sites selected based on various configurations (random, cost function, targeted submodule) for benchmarks: (a) ORCTRL and (b) FIR Filter.
Fig. 6 compares the number of target sites handled by PROFUZZ and DirectFuzz across multiple benchmarks, showing a 2× to 30× increase in the number of target sites covered by PROFUZZ. This highlights the scalability and effectiveness of our approach, demonstrating its ability to handle a much larger number of target sites, particularly for complex hardware designs. As illustrated in Fig. 7, PROFUZZ demonstrates a steady improvement in target site coverage over time for all the configurations for ORCTRL and FIR Filter, highlighting the effectiveness of its directed fuzzing and efficient seed exploration in covering target sites.